OpenStudy (swimmerranae102000):

Given that M is between A and Z. If ZM=3x+2 AM=3x+4 Az=9x-3 find X and AM

jhonyy9 (jhonyy9):

so bc. M is between A and Z from this result that AM + ZM = AZ - yes ?

jhonyy9 (jhonyy9):

rewrite this equation using these given values and solve it for x

jhonyy9 (jhonyy9):

ok. ? please collaborate

jhonyy9 (jhonyy9):

are you here please ?

OpenStudy (swimmerranae102000):

I'm sorry

jhonyy9 (jhonyy9):

you dont understand these above wrote ?

OpenStudy (swimmerranae102000):

Ya I'm working on it

OpenStudy (swimmerranae102000):

(3x+4) +(3x+2)=9x-3

jhonyy9 (jhonyy9):

so there are given AZ = 9x -3 MZ = 3x +2 AM = 3x+4 and we know that M is between A and Z so from this result that AZ = ZM + AM yes ?

OpenStudy (swimmerranae102000):

Yes

jhonyy9 (jhonyy9):

yes right so from this you need calculi the value of x do oyu can solve it for x ?

jhonyy9 (jhonyy9):

do you sorry

OpenStudy (swimmerranae102000):

I'm not to sertant how to

OpenStudy (swimmerranae102000):

Would you use distributive property

jhonyy9 (jhonyy9):

(3x+4) +(3x+2)=9x-3 you get this equation so write it without parentheses

jhonyy9 (jhonyy9):

and assume the same terms

jhonyy9 (jhonyy9):

without parentheses will see it more clearly

jhonyy9 (jhonyy9):

3x+4+3x+2=9x-3 3x+3x+4+2=9x-3 continue it please

OpenStudy (swimmerranae102000):

i gtg I have to do this later

jhonyy9 (jhonyy9):

ok. bye

OpenStudy (phi):

As jhon was saying, first add ZM=3x+2 and AM=3x+4 in other words, add 3x+2+3x+4 when you add things, you can change the order (example: 1+2 is the same as 2+1) so write it as 3x+3x+2+4 I assume you can add the last part: the 2+4. what is that ?

OpenStudy (swimmerranae102000):

wait can some one help